### v1

# Investigating Solid-Fluid Phase Coexistence in DC Plasma Bilayer Crystals: The Role of Particle Pairing and Mode Coupling

Abstract
This article presents a detailed investigation of solid-fluid phase coexistence in a bilayer dusty plasma crystal subjected to varying confinement ring bias voltages in a DC glow discharge argon plasma. Melamine formaldehyde particles were employed to form a stable, hexagonally ordered bilayer crystal within a confinement ring electrically isolated from the grounded cathode. By systematically adjusting the confinement ring bias, a distinct phase coexistence emerged: it is characterized by a fluid-like melted core surrounded by a solid crystalline periphery. Crucially, analysis of the phonon spectra revealed frequency shifts that deviate significantly from the predictions of classical monolayer Mode-Coupling Instability (MCI) theory. Stability analysis further demonstrated that dynamic interlayer particle pairing and non-reciprocal interactions play a pivotal role in destabilizing the bilayer structure. These findings highlight previously underappreciated mechanisms driving the melting transition in bilayer dusty plasmas, offering a more comprehensive understanding of phase behavior in complex plasma systems. The results underscore the importance of interlayer coupling and confinement effects in tuning structural transitions.
preprint: AIP/123-QED
I Introduction
Dusty plasma crystals, consisting of monodisperse micrometer-sized particles suspended in a gas discharge plasma, provide a unique platform to study phase transitions [1, 2, 3, 4, 5, 6], stability [7, 8], and melting dynamics in systems with long-range interactions [9]. Micron or sub-micron sized particles once inserted into the plasma, become highly charged due to interactions with the surrounding ions and electrons and form stable, lattice-like structures under suitable confinement [10, 11, 12], giving rise to what is termed a “plasma crystal.” Such systems are analogous to soft condensed matter systems and offer a controlled means of investigating particle-particle interactions, solid-fluid coexistence, and melting processes, making them valuable for studies that bridge plasma physics and condensed matter theory [13, 14].
In recent years, dusty plasma crystals have emerged as a model system for exploring phase coexistence [15, 16, 17, 18, 19, 20, 21], where solid and fluid phases can coexist within the same structure. The dynamics underlying this phenomenon are complex, as they involve particle interactions, structural rearrangements, and instabilities that are not commonly observed in simpler systems. Phonon modes play a critical role in understanding phase transitions and melting dynamics in complex plasmas. Longitudinal and transverse phonon modes, arising from compressional and shear motions respectively, provide insights into system stability and energy distribution [22, 23, 24]. Wang et al. [22] developed a theoretical framework for phonon modes in Yukawa crystals a model analogous to dusty plasmas with screened Coulomb interactions. This enabled systematic studies of dispersion relations and the influence of confinement and interaction strength on phonon behavior and instabilities  [25, 26, 7, 27].
In bilayer or multilayer complex plasma crystals, the Schweigert instability is a commonly observed phenomenon that results in the melting of the crystal structure [3, 28, 29]. It originates from the non-reciprocal nature of interactions between particles in adjacent layers due to constant ion flows, which produce positively charged regions ion wakefields downstream of each particle [30, 31, 32]. These ion wakes exert attractive forces on nearby particles, rendering the microparticle interactions non-reciprocal. In multilayer systems, ion wakes of upper-layer particles attract lower-layer particles, but no corresponding force acts in the opposite direction. This asymmetry leads to a non-reciprocal force balance that destabilizes vertical particle alignment and results in growing horizontal oscillations a process identified as the Schweigert instability that can ultimately melt the ordered plasma crystal [15]. Zhdanov et al. [33, 34] reported spontaneous particle pairing and collective dynamics in two-dimensional plasma crystals, showing that the accumulation of such pairs contributes to structural melting.
Studies on phase transitions in monolayer complex plasmas have shown that mode-coupling instability (MCI) is the primary mechanism responsible for melting.  [35, 7, 26, 34, 27]. It was observed that the coupling between longitudinal and out-of-plane modes triggers the destabilization of the lattice, ultimately leading to melting. Ivlev et al. [36] later proposed a common framework for monolayer and bilayer crystals, highlighting wake-mediated non-reciprocity, MCI and a universal stability principle for both monolayer and bilayer complex plasmas. The instability is driven by the coupling of wave modes, which is caused by nonreciprocal, wake-mediated interparticle forces. In monolayers, ions wakes introduce a weak negative coupling between the normally separate horizontal and vertical modes. Conversely, bilayers possess an inherent positive coupling between horizontal modes due to their structural asymmetry. Therefore, for a bilayer to become unstable, the wake-induced negative coupling must be strong enough to overcome this existing positive offset, a condition that is facilitated in both systems when modes are in resonance. Recently, Jaiswal et al. [16] reported the first experimental observation of phase coexistence in bilayer dust clouds, induced by confinement variations. While they observed signs of MCI, the absence of intersection between vertical and horizontal modes and phase coexistence seemed to emerge when interlayer and interparticle spacing approach or fall below the screening length. This highlights the need for detailed investigation of wake-mediated non-reciprocal interactions and their role in bilayer crystal stability.
This article investigates the mechanisms underlying phase coexistence and melting in bilayer dusty plasma crystals, with a specific focus on the interplay between particle pairing and mode coupling under varying confinement ring bias voltages in a DC glow discharge plasma. By analyzing phonon spectra across a range of confinement potentials, we explore how longitudinal and transverse modes respond to particle interactions and coupling effects. Through detailed tracking of particle trajectories, separations, and velocities, we uncover new insights into the dynamic processes that destabilize the crystal structure. Our results demonstrate that dynamic particle pairing, combined with a modified form of MCI, serves as a primary driver of phase transitions, ultimately leading to melting. Additionally, we introduce and apply a quantitative analysis of interlayer non-reciprocity, capturing the emergence of asymmetric particle interactions that signal structural instability. This approach provides a novel diagnostic for identifying the onset of disorder and melting in bilayer dusty plasmas. Overall, this study contributes to a deeper understanding of structural stability and phase transitions in complex plasma systems, offering new perspectives on the microscopic interactions that govern these phenomena.
Figure 1: Schematic of the experimental arrangement. L1 and L2 are green (532 nm) and red (650 nm) illumination lasers.
Figure 2: Snapshot showing overlapped image of 20 consecutive frames from the top view camera at 103 V of confining ring voltage. Circular marked region shows the region of melting. The outside region is almost stationary with a small thermal fluctuation around the equilibrium position.
Figure 3: View of the dust cloud from the side camera at different bias voltage. The layer gap decreases with reducing confinement potential from 140 V to 103 V.
II Experimental procedure
The experimental setup closely follows the configuration described by Jaiswal et al. [16] in the Low-Temperature Dusty Plasma Experimental device (LDPEx). A schematic of the setup is shown in Fig. 1. The gas discharge was initiated in a borosilicate glass tube measuring 33 cm in length and 7.6 cm in inner diameter. The tube is equipped with multiple axial and radial ports to enhance accessibility and diagnostic capabilities. A stainless-steel (SS) disk anode (2.5 cm diameter) and a 30 cm $\times$ 6 cm $\times$ 0.6 cm SS plate cathode were positioned 6 cm apart inside the chamber. This asymmetric electrode configuration minimizes ion heating of the dust particles, enabling the formation of a stable dust plasma crystal in the DC glow discharge environment. The chamber was initially pumped down to a base pressure of $10^{-3}$ mbar, then filled with argon gas to maintain a working pressure of 112 mTorr (15 Pa).
[…]
Figure 4: Longitudinal (top) and transverse (bottom) spectra for $0^{\circ}$ k-vector orientation
Figure 5: Longitudinal (top) and transverse (bottom) spectra for $30^{\circ}$ k-vector orientation
Figure 6: Optical phonon modes for different bias voltages
III Results
Fig. 2 shows an overlapped image of 20 consecutive frames from the top view camera at 103 V of confining ring voltage. The circular marked region shows a clear melting whereas the periphery is almost stationary with a small thermal fluctuation around the equilibrium position. The side view of the particle cloud (shown in Fig. 3) reveal that the interlayer gap between two layer is decreasing as the confinement potential decreases from 140 V to 103 V. This increases the interaction between the layers and vertical oscillations and coupling of two layer were seen for the voltage below 120 V (see supplementary material).
III.1 Phonon spectra
The phonon spectrum of the dust lattice [35, 33] was analyzed over a range of confinement ring bias voltages, focusing on optical modes (from side-view data) and acoustic modes (from top-view data). Particle positions were extracted using ImageJ [37] for both views. The extracted positions were processed using the Trackpy library [38, 39] to link particle trajectories and compute velocities. Longitudinal and transverse phonon currents were then calculated following the method of Nunomura et al. [23], and Fourier transforms were applied to obtain the current fluctuation spectra. This analysis across various ring bias voltages enabled us to study the evolution of wave modes in the crystal.
As the confinement bias decreased from 140 V to 103 V, we observed a notable increase in the maximum frequencies of both longitudinal and transverse acoustic modes, rising from 17.5 Hz to 22.5 Hz (Figs. 55). Similarly, optical mode frequencies increased from approximately 30 Hz at 140 V to around 39 Hz at 103 V (Fig. 6), a trend also visible in the normalized power spectral density [16]. This upward frequency shift suggests that reducing the ring bias voltage stiffens the confinement potential, thereby strengthening interparticle interactions and enhancing collective dynamics.
At a confinement bias of 120 V, signs of torsional motion [40] become visible, indicating the onset of particle pairing [see hotspots in Figs. 5(c), (g) and 5(c), (g)]. As previously reported, torsion and MCI develop over similar spatial and temporal scales [40]. However, in contrast to monolayer MCI, no localized hot spots are present in the out-of-plane spectra (Fig. 6(c)). Below 120 V, spectral features appear increasingly blurred. It is possible that the imaging setup captured particles from both layers, leading to artificial mixing of spectral features. This behavior likely reflects a change in interparticle coupling, possibly due to the formation of particle pairs between the two layers. Such pairing can modify the system’s vibrational response and affecting the mechanical stability of the lattice [40].
The out-of-plane spectra show an enhancement in vertical motion (Fig. 6), akin to classical monolayer MCI. [35, 7] Furthermore, traces of the out-of-plane mode becomes visible in the in-plane spectra around 35–40 Hz for voltages below 120 V, suggesting interlayer mode coupling. While some features such as acoustic mode resonances resemble traditional MCI, the absence of a corresponding optical hot spots and the presence of harmonics point to a modified instability. This bilayer-specific variant likely stems from vertical asymmetry in the interaction potential and dynamic pairing behavior. These harmonics support the view that vertical coupling and interlayer asymmetry significantly reshape phonon mode structure, resulting in modified instability mechanisms unique to bilayer complex plasma systems.
III.2 Dependence of particle pairing on the confining ring bias voltage
As the ring bias voltage was reduced, a noticeable increase in the number of particle pairs formed was observed. Tracking the number of pairs as a function of voltage revealed a clear trend: the number of pairs increased as the voltage decreased (Fig. 7), particularly around 120 V and below. At 103 V, the number of pairs was measured as 107 ± 10, confirming the strong increase in pair formation at lower confinement potentials
At higher ring bias voltages, such as 140 V, particle interactions were weaker, and pair formation occurred less frequently. As the voltage decreased, the frequency of pairing events increased, indicating stronger interactions. This observation aligns with the phonon spectrum data, where the maximum phonon frequency increased with decreasing voltage, reflecting a more strongly coupled system. The intensifying interactions at lower voltages appear to destabilize the structure, pushing the system toward a more fluid-like state.
Overall, the relationship between voltage and pair formation underscores the critical role of particle pairing in the transition from a solid-like phase to a more disordered, fluid state.
Figure 7: Evolution of the number of particle pairs observed as function of the confining ring bias voltage. Pair counts were performed manually on top-view frames; uncertainty is reported as the counting (Poisson) error $\sqrt{N}$ and conservatively covers human error in identifying overlapping/masked particles.
III.3 Case study: dynamics of particle pairing at a confining ring bias voltage of 120 V
As we have seen in previous sections, At 120 V, significant changes in particle behavior were observed. For example, the hot spot arises at 120 V and the pairing effect between the particles also increases drastically after 120 V. This has important consequence on particle structural dynamics therefore 120 V has been considered as a critical voltage for understanding the transition from crystalline to fluid states. In order to illustrate that, we focus on three particles: "1" and "2" in the top layer (Layer 1 in Fig. 8(a)), and "3" just above Layer 1. Their motion over 6 seconds reveals the dynamics of pair formation, collapse, and particle dragging, contributing to system destabilization. Fig. 8(b) shows the evolution of their $x$-coordinates.
Figure 8: (a) Coordinates of particles in the region of interest. (b) Horizontal trajectory of particles 1 (blue), 2 (green), and 3 (red). (c) Fluctuations in horizontal velocity of particles. The dashed lines represent the average horizontal velocity of each particle. (d) Horizontal separation between particles in both pairs. (e) Vertical separation between particles in both pairs.
Initially, particle "3" formed a stable pair with "1" for $\sim$1.5 s, as shown by their nearly identical $x$-coordinates. At $t\simeq 1.5$ s, the pair broke, and "3" drifted leftward toward particle "2," forming a new pair shortly before $t=2$ s. Around $t\simeq 3$ s, particle "3" continued leftward and while dragging along particle "2" which gradually moved leftwards too. The pair realigned by $t\simeq 4$ s. This dragging is the signature of the ion wake below particle "3" influences nearby particles (also evidence by Melzer et al.[30]), potentially destabilizing the entire system.
Fig. 8(c) represents the horizontal velocity fluctuations which provide a direct measure of energy exchange during the pairing process. The average horizontal velocities were $\langle V_{x_{1}}\rangle$ = 0.747 mm/s, $\langle V_{x_{2}}\rangle$ = 1.234 mm/s, and $\langle V_{x_{3}}\rangle$ = 1.298 mm/s, indicating that particle 3 exhibits the strongest horizontal motion, consistent with its role in successive pair interactions. Distinct velocity spikes in particle 3 occur during its transition from pairing with particle 1 to forming a pair with particle 2, as well as during the intermediate phase where it drags particle 2 signifying a moment of energy transfer and structural rearrangement. A corresponding peak in the velocity of particle 2 confirms their coupled motion. Once particle 3 establishes a stable pairing with particle 2, the velocities of both particles stabilize, indicating the re-establishment of a transient equilibrium. These correlated variations in $V_{x}$ highlight the dynamic exchange of momentum during pair formation and collapse, central to the onset of local disorder in the bilayer.
Horizontal separation as shown in Fig. 8(d) remained constant for the first 1.2 s, then diverged:
  • 1-3 pair: separation increased after 1.2 s, stabilized around 3.9 s.
  • 2-3 pair: separation decreased after 1.2 s, constant until 3 s, then increased at 3.6 s as "3" moved leftward. It decreased again as "2" was pulled left, stabilizing around 4.2 s.
Vertical separation [Fig. 8(e)] remained mostly constant at $0.06\pm 0.01$ mm , indicating that the dragging was primarily horizontal. These trends reflect how particle "3" influenced the motion and separation of particles "1" and "2" over time.
The horizontal dragging force of particle "3" on particle "2" was estimated as:
$F_{x}=\gamma\cdot\langle V\rangle\cdot m=0.413~\text{fN}$
where $\gamma$ is the Epstein drag coefficient [41], $\langle V\rangle$ is the average horizontal velocity and $m$ is the mass of the particles. Assuming the system is overdamped and using $F_{x}=k\cdot\Delta x$, the spring constant for the 2-3 pair during the interaction was:
$k=3.75~\text{pN}\cdot\text{m}^{-1}$
The estimated horizontal dragging force and effective spring constant quantify the strength of the wake-mediated coupling between paired particles. These values confirm that the interparticle attraction is sufficiently strong to produce measurable lateral displacement and energy exchange. The spring constant thus characterizes how mechanical energy is stored and released during pairing events, linking microscopic force asymmetry to the observed velocity fluctuations and, ultimately, to the onset of melting. [34].
The dragging of "2" by "3" provides key insight into the melting transition. Increasing interaction strength, reduced particle distance, and energy transfer from "3" disrupt the crystal structure. Continuous pair formation and collapse, such as observed between particles "2" and "3", drive the system toward a fluid-like state. Velocity fluctuations and changing horizontal separation suggest that energy transfer during pairing events is a primary driver of melting. These dynamic interactions locally destabilize the structure, leading to crystallinity breakdown.
III.4 Nonreciprocal interactions and driving strength
In complex plasmas, the effective interaction between microparticles is mediated by streaming ions. The resulting wakefields produce an inherently anisotropic potential that depends on flow direction and spatial asymmetry. In bilayer systems, this anisotropy translates into a fundamental breakdown of Newton’s third law: the attractive force experienced by a lower-layer particle due to the ion wake of an upper-layer particle is not balanced by a corresponding force in the opposite direction. This nonreciprocity makes bilayer crystals qualitatively different from equilibrium Yukawa systems, as the microscopic forces themselves carry a directional bias that drives the system away from detailed balance.
To quantify this effect, we introduced a pair-resolved metric following from the recent work from Ivlev et al. [42]
$\mathbf{R}=<|F_{top}+F_{bottom}|>$ (1)
which directly probes the imbalance of forces in paired particles across the two layers. In the case of reciprocal interactions, forces cancel and R vanishes, whereas any finite value signals that wake-mediated asymmetry injects net momentum into the system. This measure has the advantage of being experimentally accessible from trajectory data, avoiding the need for coarse-grained approximations or indirect stability proxies.
Figure 9: Non-reciprocal driving strength $R$ versus bias voltage. Shaded regions show the distribution of $R$ values, while points and error bars denote the spread of the data.
We have calculated the value of R for all particle pairs, and the results show a clear dependence on the confinement bias voltage. Fig. 9 illustrates how the distribution of R varies with changes in confinement voltage. In the figure, the shaded regions represent the spread of pairwise non-reciprocal driving strengths at each voltage. A wider shaded region indicates greater variation in R among particle pairs. As the confinement voltage decreases, the spread of the distribution increases, indicating enhanced non-reciprocity and variability. At 140 V, the measured R is small, indicating that interlayer forces are relatively reciprocal and the crystal retains long-range order. Reducing the confinement to 120 V increases interlayer overlap and enhances wakefield-mediated attraction, raising the value of R. This coincides with a marked increase in particle-pairing events. At 110 V, where the bilayer collapses into a strongly fluctuating state, the nonreciprocal driving strength reaches its maximum, confirming that asymmetric forces dominate the dynamics at low confinement. The monotonic growth of R across this range establishes a direct experimental link between microscopic nonreciprocity and macroscopic melting.
The importance of this observation is underscored by theoretical work by Ivlev et al.[42], which demonstrates that nonreciprocal interactions fundamentally alter statistical mechanics. Unlike equilibrium Yukawa crystals, where equipartition ensures a well-defined temperature, nonreciprocal systems can sustain different effective kinetic temperatures for different particle species, or even exhibit continuous energy growth with no steady-state equilibrium. Our data support the latter scenario: as non-reciprocity increases, particle pairs act as channels for energy injection from the ion flow into the lattice. The repeated formation and collapse of pairs provides the necessary energy to local rearrangements and velocity fluctuations, ultimately destabilizing the entire crystal.
An additional consequence of non-reciprocity is the emergence of a dynamical asymmetry in how instabilities develop. Classical Mode-Coupling Instability (MCI) is triggered when longitudinal and vertical phonon modes intersect, producing resonant energy transfer. In contrast, in a bilayer system, the phonon spectra exhibit MCI-like features (frequency shifts, harmonics) but without the expected resonance crossing of longitudinal and vertical modes. This indicates that the instability originates from the combined action of wake-driven non-reciprocity and modified mode coupling, rather than classical MCI alone.
In summary, the analysis of nonreciprocal interactions demonstrates that melting in bilayer dusty plasmas is not only driven by damping thresholds or classical instabilities, but also by a confinement-controlled increase in wake-mediated asymmetry. The monotonic rise of nonreciprocal driving strength with decreasing confinement voltage provides a quantitative, experimentally validated mechanism for how crystalline order gives way to fluid-like coexistence. This establishes nonreciprocity not only as a descriptive feature of bilayer plasmas but as the primary physical agent responsible for their phase transition.
IV Conclusion
In this work, we explored the mechanisms driving solid–fluid phase coexistence and melting in a bilayer dusty plasma crystal in a DC discharge by combining experimental phonon spectra, particle-level dynamics, and a new quantitative measure of non-reciprocity. Our aim was to investigate the interplay between Mode-Coupling Instability (MCI), particle pairing, and structural disorder under varying confinement conditions.
Phonon spectral analysis revealed MCI-like hotspots near 19 Hz in the acoustic modes, while vertical oscillations in the optical modes occurred in the 30–39 Hz range well above the classical MCI resonance. These results confirm that instability in bilayers differs fundamentally from monolayer MCI, being strongly influenced by interlayer coupling and structural asymmetry. The Schweigert mechanism is also not applicable here: we observed clear vertical oscillations in both layers, growing with decreasing confinement bias, even at constant neutral pressure. This indicates that melting is governed by confinement-driven structural changes rather than a damping threshold.
Side-view imaging revealed repeated formation, collapse, and exchange of particle pairs. These dynamic interactions disrupted local order, redistributed energy, and strongly influenced collective modes. To quantify their impact, we introduced a non-reciprocity metric based on pair accelerations. This analysis showed a clear monotonic increase of non-reciprocal driving strength with decreasing confinement voltage, directly linking microscopic wake-mediated asymmetry to macroscopic instability. Stronger non-reciprocity at low voltages coincided with enhanced particle pairing and eventual melting.
Together, these findings point to a hybrid mechanism for melting in bilayer dusty plasma crystals, where vertical confinement, wakefield-mediated non-reciprocal interactions, dynamic particle pairing, and local structural fluctuations act cooperatively to destabilize the crystalline order. This establishes dynamic particle pairing amplified by non-reciprocal energy injection as the key driver of the phase transition. The framework developed here, which combines phonon spectral analysis with direct quantification of non-reciprocity, provides a general approach for studying melting, energy transport, and defect dynamics in strongly coupled non-equilibrium systems.